OpenStudy (anonymous):

Percy paid $24.10 for a basketball. The price of the basketball was $22.99. What was the sales tax rate?

OpenStudy (anonymous):

yes i am in consumers mathmatics

OpenStudy (anonymous):

Paid/price tag*100 is difference in % from price tag.

OpenStudy (mayankdevnani):

4.58 is answer right @lilhen

OpenStudy (mayankdevnani):

or 4.8

OpenStudy (anonymous):

yes indeed

OpenStudy (mayankdevnani):

4.8 is answer

OpenStudy (anonymous):

ok

OpenStudy (mayankdevnani):

To figure out what the sales tax rate is, we use a simple formula: Original Price plus a certain percentage of the original price equals the sales price. Written mathetmatically: x + %x = y x = original price y = sales price % = sales rate In this case: x = 22.99 y = 24.10 % = unknown So inputting all the known values into the formula, we get: 22.99 + %(22.99) = 24.10 Now we solve for %, which is the sales rate: 22.99 + %(22.99) = 24.10 Subtract 22.99 from both sides: %(22.99) = 24.10 - 22.99 %(22.99) = 1.11 Divide both sides by 22.99: % = 1.11/22.99 % = 4.8

OpenStudy (mayankdevnani):

ok @lilhen

OpenStudy (mayankdevnani):

am i right @jiteshmeghwal9

OpenStudy (anonymous):

yes thank you tht was simpler than how im being taught ill save this

OpenStudy (mayankdevnani):

ok welcome

OpenStudy (jiteshmeghwal9):

u r really a commercial mathematics teacher :)

OpenStudy (mayankdevnani):

thnx.. @jiteshmeghwal9